OpenStudy (anonymous):

how many solutions does this eqn have? |3x+7|=0 please explain.

OpenStudy (anonymous):

only way for the absolute value of something to be zero is if the expression inside is 0 how many solutions are there to \(3x+7=0\) ?

OpenStudy (anonymous):

sorry i wrote the problem wrong. hold on

OpenStudy (anonymous):

|3x+7|+2=0

hero (hero):

Solve it and see

hero (hero):

If you subtract 2 from both sides you get |3x + 7| = -2 Absolute value means distance and you can have positive distance or no distance, but what you can't have is negative distance.